# A Bias–Variance Decomposition for Ensembles over Multiple Synthetic Datasets

### 06 Feb 2024 v1

@@ Abstract
Recent studies have highlighted the benefits of generating multiple synthetic datasets for supervised learning, from increased accuracy to more effective model selection and uncertainty estimation. These benefits have clear empirical support, but the theoretical understanding of them is currently very light. We seek to increase the theoretical understanding by deriving bias-variance decompositions for several settings of using multiple synthetic datasets. Our theory predicts multiple synthetic datasets to be especially beneficial for high-variance downstream predictors, and yields a simple rule of thumb to select the appropriate number of synthetic datasets in the case of mean-squared error and Brier score. We investigate how our theory works in practice by evaluating the performance of an ensemble over many synthetic datasets for several real datasets and downstream predictors. The results follow our theory, showing that our insights are also practically relevant.
@@ 1 Introduction
Synthetic data has recently attracted significant attention for several application in machine learning. The idea is to generate a dataset that preserves the population-level attributes of the real data, making the synthetic data useful for analysis, while also accomplishing a secondary task, such as improving model evaluation (van Breugel et al., 2023b), fairness (van Breugel et al., 2021), data augmentation (Antoniou et al., 2018; Das et al., 2022) or privacy (Liew et al., 1985; Rubin, 1993). For the last application, differential privacy (DP) (Dwork et al., 2006) is often combined with synthetic data generation (Hardt et al., 2012; McKenna et al., 2021) to achieve provable privacy protection, since releasing synthetic data without DP can be vulnerable to membership inference attacks (van Breugel et al., 2023c; Meeus et al., 2023).
Several lines of work have considered generating multiple synthetic datasets from one real dataset for various purposes, including statistical inference (Rubin, 1993; Raghunathan et al., 2003; Räisä et al., 2023a) and supervised learning (van Breugel et al., 2023a). We focus on the latter setting.
@@
However, van Breugel et al. (2023a) have very little theoretical analysis of the generative ensemble. They justify the ensemble by assuming the synthetic data is generated from a posterior predictive distribution. This assumption can be justified heuristically for deep generative models through a Bayesian interpretation of deep ensembles (Lakshminarayanan et al., 2017; Wilson and Izmailov, 2020; Wilson, 2021). However this justification only applies to generators with highly multi-modal losses, like neural networks. It also does not provide any insight on how different choices in the setup, like the choice of downstream predictor, affect the performance of the ensemble.
The bias-variance decomposition (Geman et al., 1992) and its generalisations (Ueda and Nakano, 1996; Wood et al., 2023) are classical tools that provide insight into supervised learning problems. The standard bias-variance decomposition from Geman et al. (1992) considers predicting $y\in\mathbb{R}$ given features $x\in\mathcal{X}$, using predictor $g(x;D)$ that receives training data $D$. The mean-squared error (MSE) of $g$ can be decomposed into bias, variance, and noise terms:
$\underbrace{\operatorname*{\mathbb{E}}_{D,y}[(y-g)^{2}]}_{\mathrm{MSE}}=\underbrace{(f(x)-\operatorname*{\mathbb{E}}_{D}[g])^{2}}_{\mathrm{Bias}}+\underbrace{\operatorname*{Var}_{D}[g]}_{\mathrm{Variance}}+\underbrace{\operatorname*{Var}_{y}[y]}_{\mathrm{Noise}}$ (1)
where we have shortened $g=g(x;D)$ and $f(x)=\operatorname*{\mathbb{E}}_{y}[y]$. $x$ is considered fixed, so all the random quantities in the decomposition are implicitly conditioned on $x$. While MSE is typically only used with regression problems, the decomposition also applies to the Brier score (Brier, 1950) in classification, which is simply the MSE of class probabilities.
@@
    We derive a bias-variance decomposition for the MSE or Brier score of generative ensembles in Theorem 2.2. This decomposition is simply a sum of interpretable terms, which makes it possible to predict how various choices affect the error. In particular, the number of synthetic datasets only affects the variance-related terms in the decomposition, predicting that generative ensembles work best with high-variance and low bias models, like interpolating decision trees.
@@
    We derive a simple rule of thumb from our decomposition to select the number of synthetic datasets in Section 2.3. In summary, 2 synthetic datasets give 50% of the potential benefit from multiple synthetic datasets, 10 give 90% of the benefit and 100 give 99% of the benefit. This also applies to bagging ensembles like random forests, which is likely to be of independent interest.
@@
    We generalise the decomposition of Theorem 2.2 to differentially private (DP) generation algorithms that do not split their privacy budget between the multiple synthetic datasets111 Theorem 2.2 applies to DP algorithms that split the privacy budget, but it is not clear if multiple synthetic datasets are beneficial with these algorithms, as splitting the privacy budget between more synthetic datasets means that each one requires adding more noise, degrading the quality of the synthetic data. in Theorem 2.4.
@@ 1.1 Related Work
Ueda and Nakano (1996) were the first to study the MSE bias-variance decomposition with ensembles, and Gupta et al. (2022); Wood et al. (2023) later extended the ensemble decomposition to other losses. All of these also apply to generative ensembles, but they only provide limited insight for them, as they do not separate the synthetic data generation-related terms from the downstream-related terms.
@@ 2 Bias-Variance Decompositions for Generative Ensembles
In this section, we make our main theoretical contributions. We start by defining our setting in Section 2.1. Then we derive a bias-variance decomposition for generative ensembles in Section 2.2 and derive a simple rule of thumb that can be used to select the number of synthetic datasets in Section 2.3. The generative ensemble decomposition doesn’t apply to several differentially private synthetic data generation methods that generate all synthetic datasets conditional on a single privatised value, so we generalise the decomposition to apply to those in Section 2.4. We also present a third decomposition that applies to Bregman divergences in Appendix B.
@@ 2.1 Problem Setting
We assume the synthetic datasets are generated by first i.i.d. sampling parameters $\theta_{1:m}|D_{r}$ given the real data, and then sampling synthetic datasets $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ given the parameters. This is how synthetic datasets are sampled in DGE (van Breugel et al., 2023a), where $\theta_{i}$ are the parameters of the generative neural network from independent training runs. This also encompasses bootstrapping, where $\theta_{i}$ would be the real dataset [^2] The $\theta_{i}$ random variables are i.i.d. if they are deterministically equal., and $p(D_{s}|\theta_{i})$ is the bootstrapping.
We will also consider a more general setting that applies to some differentially private synthetic data generators that do not fit into this setting in Section 2.4.
@@ 2.2 Mean-squared Error Decomposition
[] Let generator parameters $\theta_{1:m}$ be i.i.d. given the real data $D_{r}$, let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ independently, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i})$. Then the mean-squared error in predicting $y$ from $x$ decomposes into six terms: model variance (MV), synthetic data variance (SDV), real data variance (RDV), synthetic data bias (SDB), model bias (MB), and noise $\operatorname*{Var}_{y}[y]$:
@@
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data, $\theta$ is a single sample from the distribution of $\theta_{i}$, $D_{s}\sim p(D_{s}|\theta)$ is a single sample of the synthetic data generating process, and $f_{\theta}$ is the optimal predictor for the synthetic data generating process given parameters $\theta$. All random quantities are implicitly conditioned on $x$. The proofs of all theorems are in Appendix A.
We can immediately make actionable observations:
@@
    Increasing the number of synthetic datasets $m$ reduces the impact of MV and SDV. This means that high-variance models, like interpolating decision trees or 1-nearest neighbours, benefit the most from multiple synthetic datasets. These models should be favoured over low variance high bias models, even if the latter were better on the real data.
@@
We can also consider two extreme scenarios:
@@
    $\mathrm{MV}$ $=\operatorname*{Var}_{D_{r}}[g(x;D_{r})]$ (10)
    $\mathrm{SDV}$ $=0$ (11)
    $\mathrm{RDV}$ $=0$ (12)
    $\mathrm{SDB}$ $=0$ (13)
    $\mathrm{MB}$ $=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]$ (14)
@@
    Return the real data: $D_{s}|D_{r}=D_{r}$ deterministically333 Theorem 2.2 applies in this scenario even with multiple synthetic datasets, as the deterministically identical synthetic datasets are independent as random variables. .
    $\mathrm{MV}$ $=0$ (15)
    $\mathrm{SDV}$ $=0$ (16)
    $\mathrm{RDV}$ $=\operatorname*{Var}_{D_{r}}[g(x;D_{r})]$ (17)
    $\mathrm{SDB}$ $=0$ (18)
    $\mathrm{MB}$ $=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]$ (19)
@@
While both of these scenarios are unrealistic, they may be approximated by a well-performing algorithm. A generator that is a good approximation to either scenario could be considered to generate high-quality synthetic data. Multiple synthetic datasets are only beneficial in the first scenario, which shows that the synthetic data generation algorithm should have similar randomness to the real data generating process, and not just return a dataset that is close the the real one with little randomness.
@@ 2.3 Estimating the effect of Multiple Synthetic Datasets
Since MSE depends on $m$, we can add a subscript to denote the $m$ in question: $\mathrm{MSE}_{m}$. Now (20) gives
@@
If we have two synthetic datasets, we can estimate $\mathrm{MV}+\mathrm{SDV}$:
$\begin{split}\mathrm{MV}+\mathrm{SDV}=2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2}).\end{split}$ (22)
@@ 2.4 Differentially Private Synthetic Data Generators
Generating and releasing multiple synthetic datasets could increase the associated disclosure risk. One solution to this is differential privacy (DP) (Dwork et al., 2006; Dwork and Roth, 2014), which is a property of an algorithm that formally bounds to privacy leakage that can result from releasing the output of that algorithm. DP gives a quantitative upper bound on the privacy leakage, which is known as the privacy budget. Achieving DP requires adding extra noise to some point in the algorithm, lowering the utility of the result.
If the synthetic data is to be generated with DP, there are two possible ways to handle the required noise addition. The first is splitting the privacy budget across the $m$ synthetic datasets, and run the DP generation algorithm separately $m$ times. Theorem 2.2 applies in this setting. However, it is not clear if multiple synthetic datasets are beneficial in this case, as splitting the privacy budget requires adding more noise to each synthetic dataset. This also means that the rule of thumb from Section 2.3 will not apply. Most DP synthetic data generation algorithm would fall into this category (Aydore et al., 2021; Chen et al., 2020; Harder et al., 2021; Hardt et al., 2012; Liu et al., 2021; McKenna et al., 2019, 2021) if used to generate multiple synthetic datasets.
@@
The synthetic datasets are not i.i.d. given the real data in the second setting, so the setting described in Section 2.1 and assumed in Theorem 2.2 does not apply. However, the synthetic datasets are i.i.d. given the noisy summary $\tilde{s}$, so we obtain a similar decomposition as before. {theorem}[] Let generator parameters $\theta_{1:m}$ be i.i.d. given a DP summary $\tilde{s}$ of the real data $D_{r}$, let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ independently, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i})$. Then
$\begin{split}\mathrm{MSE}&=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\mathrm{RDV}+\mathrm{DPVAR}\\ &+\,(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y],\end{split}$ (24)
@@
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data, $\theta$ is a single sample from the distribution of $\theta_{i}$, $D_{s}\sim p(D_{s}|\theta)$ is a single sample of the synthetic data generating process, and $f_{\theta}$ is the optimal predictor for the synthetic data generating process given parameters $\theta$. All random quantities are implicitly conditioned on $x$.

### 23 May 2024 v2

@@ Abstract
Recent studies have highlighted the benefits of generating multiple synthetic datasets for supervised learning, from increased accuracy to more effective model selection and uncertainty estimation. These benefits have clear empirical support, but the theoretical understanding of them is currently very light. We seek to increase the theoretical understanding by deriving bias-variance decompositions for several settings of using multiple synthetic datasets, including differentially private synthetic data. Our theory predicts multiple synthetic datasets to be especially beneficial for high-variance downstream predictors, and yields a simple rule of thumb to select the appropriate number of synthetic datasets in the case of mean-squared error and Brier score. We investigate how our theory works in practice by evaluating the performance of an ensemble over many synthetic datasets for several real datasets and downstream predictors. The results follow our theory, showing that our insights are practically relevant.
@@ 1 Introduction
Synthetic data has recently attracted significant attention for several applications in machine learning. The idea is to generate a dataset that preserves the population-level attributes of the real data. This makes the synthetic data useful for analysis, while also accomplishing a secondary task, such as improving model evaluation (van Breugel et al., 2023b), fairness (van Breugel et al., 2021), data augmentation (Antoniou et al., 2018; Das et al., 2022) or privacy (Liew et al., 1985; Rubin, 1993). For privacy, differential privacy (DP) (Dwork et al., 2006) is often combined with synthetic data generation (Hardt et al., 2012; McKenna et al., 2021) to achieve provable privacy protection, since releasing synthetic data without DP can be vulnerable to membership inference attacks (Stadler et al., 2022; van Breugel et al., 2023c; Meeus et al., 2023).
Several lines of work have considered generating multiple synthetic datasets from one real dataset for various purposes, including statistical inference (Rubin, 1993; Raghunathan et al., 2003; Räisä et al., 2023a) and supervised learning (van Breugel et al., 2023a), the latter of which is our focus.
@@
However, van Breugel et al. (2023a) have very little theoretical analysis of the generative ensemble. Their theoretical justification assumes the data is generated from a posterior predictive distribution. This assumption can be justified heuristically for deep generative models through a Bayesian interpretation of deep ensembles (Lakshminarayanan et al., 2017; Wilson & Izmailov, 2020; Wilson, 2021). However, this justification only applies to generators with highly multi-modal losses, like neural networks. It also does not provide any insight on how different choices in the setup, like the choice of downstream predictor, affect the performance of the ensemble.
The bias-variance decomposition (Geman et al., 1992) and its generalisations (Ueda & Nakano, 1996; Wood et al., 2023) are classical tools that provide insight into supervised learning problems. The standard bias-variance decomposition from Geman et al. (1992) considers predicting $y\in\mathbb{R}$ given features $x\in\mathcal{X}$, using predictor $g(x;D)$ that receives training data $D$. The mean-squared error (MSE) of $g$ can be decomposed into bias, variance, and noise terms:
$\underbrace{\operatorname*{\mathbb{E}}_{D,y}[(y-g(x;D))^{2}]}_{\mathrm{MSE}}=\underbrace{(f(x)-\operatorname*{\mathbb{E}}_{D}[g(x;D)])^{2}}_{\mathrm{Bias}}+\underbrace{\operatorname*{Var}_{D}[g(x;D)]}_{\mathrm{Variance}}+\underbrace{\operatorname*{Var}_{y}[y]}_{\mathrm{Noise}}$ (1)
where we $f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor. $x$ is considered fixed, so all the random quantities in the decomposition are implicitly conditioned on $x$. While MSE is typically only used with regression problems, the decomposition also applies to the Brier score (Brier, 1950) in classification, which is simply the MSE of class probabilities.
@@
    We derive a bias-variance decomposition for the MSE or Brier score of generative ensembles under an i.i.d. assumption in Theorem 2.2. This decomposition is simply a sum of interpretable terms, which makes it possible to predict how various choices affect the error. In particular, the number of synthetic datasets only affects the variance-related terms in the decomposition, predicting that generative ensembles work best with high-variance and low bias models, like interpolating decision trees.
@@
    We derive several practical considerations from our decomposition, including a simple rule of thumb to select the number of synthetic datasets in Section 2.3. In summary, 2 synthetic datasets give 50% of the potential benefit from multiple synthetic datasets, 10 give 90% of the benefit and 100 give 99% of the benefit. This also applies to bagging ensembles like random forests, which is likely to be of independent interest.
@@
    We generalise the decomposition of Theorem 2.2 to differentially private (DP) generation algorithms that do not split their privacy budget between the multiple synthetic datasets111 Theorem 2.2 applies to DP algorithms that split the privacy budget, but it is not clear if multiple synthetic datasets are beneficial with these algorithms, as splitting the privacy budget between more synthetic datasets means that each one requires adding more noise, degrading the quality of the synthetic data. in Theorem 2.4, to non-i.i.d. synthetic data in Appendix B, and to Bregman divergences in Appendix C.
@@ 1.1 Related Work
In a similar manner, multiple imputation (Rubin, 1987; Murray, 2018) imputes missing values multiple times to a dataset, which allows correcting uncertainty estimates for the extra uncertainty from missing values. Multiple imputation is typically used with statistical inference in mind, but the imputed datasets could be used to train an ensemble.
@@
Ueda & Nakano (1996) were the first to study the MSE bias-variance decomposition with ensembles, and Gupta et al. (2022); Wood et al. (2023) later extended the ensemble decomposition to other losses. All of these also apply to generative ensembles, but they only provide limited insight for them, as they do not separate the synthetic data generation-related terms from the downstream-related terms.
@@ 2 Bias-Variance Decompositions for Generative Ensembles
In this section, we make our main theoretical contributions. We start by defining our setting in Section 2.1. Then we derive a bias-variance decomposition for generative ensembles in Section 2.2 and derive a simple rule of thumb that can be used to select the number of synthetic datasets in Section 2.3. The generative ensemble decomposition does not apply to several differentially private synthetic data generation methods that generate all synthetic datasets conditional on a single privatised value, so we generalise the decomposition to apply to those in Section 2.4. We also present a third decomposition that applies to Bregman divergences, but is not as informative as the others, in Appendix C.
@@ 2.1 Problem Setting
We assume that the synthetic datasets are generated by first i.i.d. sampling parameters $\theta_{1:m}|D_{r}$ given the real data, and then sampling synthetic datasets $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ given the parameters. This is how synthetic datasets are sampled in DGE (van Breugel et al., 2023a), where $\theta_{i}$ are the parameters of the generative neural network from independent training runs. This also encompasses bootstrapping, where $\theta_{i}$ would be the real dataset [^2] The $\theta_{i}$ random variables are i.i.d. if they are deterministically equal., and $p(D_{s}|\theta_{i})$ is the bootstrapping.
We will also consider a more general setting that applies to some differentially private synthetic data generators that do not fit into this setting in Section 2.4. In Appendix B, we consider settings without the i.i.d. assumption.
@@ 2.2 Mean-squared Error Decomposition
[] Let the parameters for $m$ generators $\theta_{i}\sim p(\theta|D_{r})$, $i=1,\dots,m$, be i.i.d. Let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ independently, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i})$. Then the mean-squared error in predicting $y$ from $x$ decomposes into six terms: model variance (MV), synthetic data variance (SDV), real data variance (RDV), synthetic data bias (SDB), model bias (MB), and noise $\operatorname*{Var}_{y}[y]$:
@@
$\begin{split}\mathrm{MSE}&=\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]\\ \mathrm{MV}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{SDV}&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\theta|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]\end{split}\quad\quad\begin{split}\mathrm{RDV}&=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})]\\ \mathrm{SDB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|D_{r}}[f_{\theta}(x)]]\\ \mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]].\end{split}$ (4)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data, $\theta\sim p(\theta|D_{r})$ is a single sample from the distribution of the generator parameters, $D_{s}\sim p(D_{s}|\theta)$ is a single sample of the synthetic data generating process, and $f_{\theta}$ is the optimal predictor for the synthetic data generating process with parameters $\theta$. All random quantities are implicitly conditioned on $x$. The proofs of all theorems are in Appendix A.
We can derive the following practical considerations from Theorem 2.2:
@@
    Increasing the number of synthetic datasets $m$ reduces the impact of MV and SDV. This means that high-variance models, like interpolating decision trees or 1-nearest neighbours, benefit the most from multiple synthetic datasets.
  3. 3.
    If $\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]$ is invariant to the size of the synthetic dataset, the only term that can change with the size of the synthetic dataset is MV. This means that the possible benefit of generating a single large synthetic dataset is smaller than generating many synthetic datasets, which can also reduce the impact of SDV.
  4. 4.
    The quality metrics for a synthetic data generator should include metrics that compare the distribution of the synthetic data to the real distribution, instead of just comparing a synthetic dataset to the real dataset. An example of the former is the downstream performance of predictive models, and examples of the latter are comparisons between the marginals of the real and synthetic datasets.
The last point can be derived by considering two extreme scenarios:
@@
    $\begin{split}\mathrm{MV}&=\operatorname*{Var}_{D_{r}}[g(x;D_{r})]\\ \mathrm{SDV}&=\mathrm{RDV}=\mathrm{SDB}=0.\end{split}\quad\quad\begin{split}\mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]\\ \phantom{x}\end{split}$ (5)
@@
    Return the real data: $D_{s}|D_{r}=D_{r}$ deterministically.333 Theorem 2.2 applies in this scenario even with multiple synthetic datasets, as the deterministically identical synthetic datasets are independent as random variables.
    $\begin{split}\mathrm{RDV}&=\operatorname*{Var}_{D_{r}}[g(x;D_{r})]\\ \mathrm{MV}&=\mathrm{SDV}=\mathrm{SDB}=0.\end{split}\quad\quad\begin{split}\mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]\\ \phantom{x}\end{split}$ (6)
@@
While both of these scenarios are unrealistic, they may be approximated by a well-performing algorithm. The generator from Scenario 2 is the optimal generator for metrics that compare the synthetic dataset to the real dataset, while the generator of Scenario 1 is optimal for metrics that compare the synthetic data and real data distributions. Multiple synthetic datasets are only beneficial in Scenario 1, which means that metrics comparing the distributions are more meaningful when multiple synthetic datasets are considered.
@@ 2.3 Estimating the effect of Multiple Synthetic Datasets
Since MSE depends on $m$, we can add a subscript to denote the $m$ in question: $\mathrm{MSE}_{m}$. Now (7) leads to the following corollary. {corollary} In the setting of Theorem 2.2, we have
@@
Proof.
The claim follows by expanding $\mathrm{MSE}_{1}$ and $\mathrm{MSE}_{m}$ with (7). ∎
If we have two synthetic datasets, we can estimate $\mathrm{MV}+\mathrm{SDV}=2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2})$, which gives the estimator
$\mathrm{MSE}_{m}=\mathrm{MSE}_{1}-2\left(1-\frac{1}{m}\right)(\mathrm{MSE}_{1}-\mathrm{MSE}_{2}).$ (9)
@@
From (8), we see that $\mathrm{MV}+\mathrm{SDV}$ is the maximum decrease in MSE that can be obtained from multiple synthetic datasets. This means that $2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2})$, or the linear regression estimate from (10), can be used as a diagnostic to check whether generating multiple synthetic datasets is worthwhile.
@@ 2.4 Differentially Private Synthetic Data Generators
Generating and releasing multiple synthetic datasets could increase the associated disclosure risk. One solution to this is differential privacy (DP) (Dwork et al., 2006; Dwork & Roth, 2014), which is a property of an algorithm that formally bounds the privacy leakage that can result from releasing the output of that algorithm. DP gives a quantitative upper bound on the privacy leakage, which is known as the privacy budget. Achieving DP requires adding extra noise to some point in the algorithm, lowering the utility of the result.
If the synthetic data is to be generated with DP, there are two possible ways to handle the required noise addition. The first is splitting the privacy budget across the $m$ synthetic datasets, and run the DP generation algorithm separately $m$ times. Theorem 2.2 applies in this setting. However, it is not clear if multiple synthetic datasets are beneficial in this case, as splitting the privacy budget requires adding more noise to each synthetic dataset. This also means that the rule of thumb from Section 2.3 will not apply. Most DP synthetic data generation algorithms would fall into this category (Aydore et al., 2021; Chen et al., 2020; Harder et al., 2021; Hardt et al., 2012; Liu et al., 2021; McKenna et al., 2019, 2021) if used to generate multiple synthetic datasets.
@@
The synthetic datasets are not i.i.d. given the real data in the second setting, so the setting described in Section 2.1 and assumed in Theorem 2.2 does not apply. However, the synthetic datasets are i.i.d. given the noisy summary $\tilde{s}$, so we obtain a similar decomposition as before. {theorem}[] Let the parameters for $m$ generators $\theta_{i}\sim p(\theta|\tilde{s})$, $i=1,\dotsc,m$, be i.i.d. given a DP summary $\tilde{s}$. Let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ independently, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i})$. Then
$\begin{split}\mathrm{MSE}&=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\mathrm{RDV}+\mathrm{DPVAR}+\,(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y],\end{split}$ (11)
@@
$\begin{split}\mathrm{MSE}&=\operatorname*{\mathbb{E}}_{y,D_{r},\tilde{s},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]\\ \mathrm{MV}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{SDV}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\operatorname*{Var}_{\theta|\tilde{s}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{RDV}&=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})].\end{split}\quad\quad\begin{split}\mathrm{DPVAR}&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]\\ \mathrm{SDB}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|\tilde{s}}[f_{\theta}(x)]]\\ \mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]]\\ \phantom{x}\end{split}$ (12)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data, $\theta\sim p(\theta|D_{r})$ is a single sample from the distribution of the generator parameters, $D_{s}\sim p(D_{s}|\theta)$ is a single sample of the synthetic data generating process, and $f_{\theta}$ is the optimal predictor for the synthetic data generating process with parameters $\theta$. All random quantities are implicitly conditioned on $x$.